A particle moving at 30 ms-1 west has a momentum of 1598kg.ms-1 west, calculate the mass of the particle
1
Expert's answer
2019-02-13T14:39:43-0500

By the definition of the momentum we have:

"p = mv,"

here, "p" is the momentum of the particle, "m" is the mass of the particle and "v" is the velocity of the particle.

Then, from this formula, we can find the mass of the particle:

"m = \dfrac{p}{v} = \dfrac{1598 kg \cdot ms^{-1}}{30 ms^{-1}} = 53.3 kg."

Answer:

"m = 53.3 kg."